I have tried to solve this equation but I'm not sure if I'm correct.
4x+38=2.5x+45.5
What I did:
4x-2.5x=1.5x
1.5x+38=45.5
38-38 and 25.5-38
1.5x=7.5
1.5x/1.5 = 7.5/1.5
x=5
Is this right? Please show me the correct way to solve it if I'm wrong. Thanks!

Found 2 solutions by jim_thompson5910, ikleyn:
Answer by jim_thompson5910(35256)


You have the correct answer. As a way to check the answer, you can plug in the value of x and simplify both sides

4x+38=2.5x+45.5
4*(x)+38=2.5*(x)+45.5
4*(5)+38=2.5*(5)+45.5 replace every x with 5
20+38=12.5+45.5
58=58

Because we get the same result on both sides, this confirms the answer.

Answer by ikleyn(52798)

.
You can learn on how to solve a single linear equation reading the lesson
    - HOW TO solve a linear equation
in this site.